Show graphically that which of the following is one to one function
(1)f(x)=ln(x)
(2)g(x)=e^x
(3) h(x)=x³
1
Expert's answer
2020-11-29T17:55:23-0500

If some horizontal line intersects the graph of the function more than once, then the function is not one-to-one. Otherwise, the function is one-to-one.

1) "f(x)=ln(x)"



There is no horizontal line that intersects the graph of the function more than once. Therefore "f(x)=ln(x)" is one-to-one function.

2) "g(x)=e^x"



There is no horizontal line that intersects the graph of the function more than once. Therefore "g(x)=e^x" is one-to-one function.

3) "h(x)=x^3"




There is no horizontal line that intersects the graph of the function more than once. Therefore "h(x)=x^3" is one-to-one function.


Answer: all three functions are one-to-one.